If a/b x c/d=1, what is the value of ac-bd?
please include explanations!!

Answer by jim_thompson5910(35256)

(a/b)(c/d) = 1 ... Start with the given equation

a/b = 1(d/c) ... Multiply both sides by d/c.

a/b = d/c ... Multiply

a/b - d/c = 0 ... Subtract d/c from both sides.

(a/b)(c/c) - (d/c)(b/b) = 0 ... Multiply the first fraction by c/c and the second by b/b

ac/bc - bd/bc = 0 ... Multiply

(ac - bd)/bc = 0 ... Combine the fractions.

ac - bd = 0*bc ... Multiply both sides by bc.

ac - bd = 0 ... Multiply


So the answer is ac - bd = 0

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------